OpenStudy (anonymous):

Jill began to solve an equation. Which expression below shows the next step Jill needs to take

OpenStudy (gibbs):

?

OpenStudy (anonymous):

I am posting the rest now

OpenStudy (gibbs):

x = y. Then x2 = xy. Subtract the same thing from both sides: x2 - y2 = xy - y2. Dividing by (x-y), obtain x + y = y. Since x = y, we see that 2 y = y. Thus 2 = 1, since we started with y nonzero. Subtracting 1 from both sides, 1 = 0

OpenStudy (leader):

Well you are multiplying by 2 on both sides

OpenStudy (leader):

So it would have to be A.

OpenStudy (confluxepic):

True.

OpenStudy (confluxepic):

@Leader

OpenStudy (leader):

What is that supposed to mean @gibbs ??

OpenStudy (gibbs):

idk

OpenStudy (leader):

So A is your answer @humphriescody

OpenStudy (anonymous):

thank you all

OpenStudy (confluxepic):

You're welcome.